The hot-water faucet takes four minutes to fill a tub, and the cold-water faucet takes three minutes for the same job. How long will it take to fill the tub if both faucets are used?
Answer by richwmiller(17219)

1/4+1/3=1/x
3/12+4/12=1/x
7/12=1/x
12/7=x